The xtab package^{∗}

Peter Wilson Catholic University of America^{†}

             Now at peter.r.wilson@boeing.com

2000/04/09

Abstract The xtab package enables long tables to be automatically broken at page boundaries. It is an extension of the supertabular package and also reduces or eliminates some of its weaknesses.

Contents

1 Introduction 1

2 The xtab package 2 2.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 The implementation 7

List of Tables

1 The principal xtab package commands . . . . . . . . . . . . . . . . 2

1 Introduction

Although the xtab package was originally developed as part of a suite for typesetting ISO international standards [Wil96], it is also applicable for use with the L^{A}TEX standard classes. The package is an extension of the supertabular package developed by Johannes Braams and Theo Jurriens.^{1} It reduces some of the weaknesses noted in the supertabular documentation and provides additional functionality.

Section 2 provides the user manual for the package which enables long tables to be automatically broken across multiple pages. Section 3 describes the implementation.

This manual is typeset according to the conventions of the L^{A}TEX docstrip utility which enables the automatic extraction of the L^{A}TEX macro source files [GMS94].

∗This file has version number v2.3, last revised 2000/04/09.

†This work was originally performed as a Guest Researcher at the National Institute of Standards and Technology.

1supertabular.sty, version 4.1c, 7 November 1997.

2 The xtab package

The supertabular package provides for the automatic breaking of a long table across page boundaries. The extension provided here enables the heading on the table on the last page to differ from those on earlier pages of the table. The downside of the extension is that L^{A}TEX has to be run twice if the document contains a supertabular. However, L^{A}TEX is usually run at least twice for any but the simplest document in order to get cross-references and Table of Contents, etc., resolved correctly.

The current version of the extension also either cures or reduces following weaknesses in the supertabular package.^{2}

1. Sometimes the top caption of a supertabular is printed on one page and the body is printed on the following page(s). That is, there is a lonely caption.

2. Sometimes the last page of a supertabular consists of an empty table. That is, just the head and foot of the table are printed.

3. If the number of lines in the first header for the table differs from the number of lines in subsequent headers, then the continuation pages of the table may be too short or, more troubling, too long.

The weaknesses are caused by trying to guess where TEX will put a page break. The package has to guesstimate how long the next entry will be in the table and, if it is too long for the available space, it puts in its own page break. If its guess is off too much in one direction, TEX will break the page unexpectadly; if it’s off in the other direction supertabular will put in an unnecessary page break.

The xtab package has reduced, but perhaps not entirely eliminated, these weaknesses. Some hand tuning may still be required.

The principal commands available are given in Table 1.

Table 1: The principal xtab package commands

Command  Effect

\begin{xtabular}{...} This is equivalent to the normal

                             \begin{tabular}{...} environment.

You supply the specification of the columns just as for the normal tabular environment. All commands that can be used within a tabular environment can also be used within the xtabular environment. Unlike the tabular environment which prevents page breaking within the tabular, the xtabular allows page breaking, so that tabulars can extend automatically across several pages.

Continued on next page

2I have corresponded with the authors of supertabular about these.

Table 1 – continued from previous page

Command  Effect

xtabular starts off with a tabular environment and checks the amount of space left on the page as it adds each row to the tabulation. If the space left on the page is too short for another row, then it ends the current tabular, performs a page break and starts another tabular on the following page. This process is repeated until all the rows have been output. There are special commands for captioning an xtabular as a table, and also elements can be automatically inserted after each (internal) \begin{tabular} and immediately before each \end{tabular}. Do not put a xtabular in a table environment, as the table environment keeps its contents on a single page (presumably you are using xtabular because its contents are longer than one page).

\end{xtabular}  End the xtabular environment.
\begin{mpxtabular} Like the xtabular environment except

that each ‘page’ is put into a minipage first. Thus it is possible to have footnotes inside an mpxtabular. The footnote text is printed at the end of each page.

\end{mpxtabular}  End the mpxtabular environment.

Note: If any of the following commands are used, then they should be placed before the particular xtabular environment that they apply to.

\topcaption{...} A command to provide a caption for the table. The caption is placed at the top of the table.

\bottomcaption{...} A command to provide a caption for the table. The caption is placed at the bottom of the table.

Continued on next page

Table 1 – continued from previous page

Command  Effect

\tablecaption{...} A command to provide a caption for the table. The caption is placed at the default position, which is at the top of the table. Notes: You cannot use the \caption command but you can put a label after any of these captioning commands. If you want captioning, the command must be specified before the start of the supertabular environment. The \...caption{} command(s) remain in effect until changed by another

\...caption command.

\tablefirsthead{...} Defines the contents of the first occurence of the tabular head. The tabular head is some special treatment of the first row in the table. This command is optional. If used, the header must be closed by the end of line command for tabulars (e.g.,

\\).

\tablehead{...} Defines the contents of the table head on subsequent pages. For example, you might want to note that this is a continuation of the table on the previous page, as well as repeating any column headings that were given at the start of the xtabular by

\tablefirsthead.

The header must be closed like the

\tablefirsthead command.

\tablelasthead{...} Defines the contents of the table head on the last page of the table. For example, you might want to note that the table is concluded on this page. The header must be closed like the

\tablefirsthead command.
\notablelasthead  Switches off the last \tablelasthead.

A \tablelasthead stays in effect until overwritten by a new \tablelasthead or cancelled by this command.

Continued on next page

Table 1 – concluded from previous page

Command  Effect

\tabletail{...} The contents of this command are inserted before the (internal) \end{tabular} on each page except for the last page of the table. For example, you might want to note that the table is continued on the next page.

\tablelasttail{...} The contents of this command are inserted before the final (internal)

                           \end{tabular} of the table.

For example, you might want to note that this is where the table ends.

As well as the xtabular and mpxtabular environments there are the corresponding starred versions (i.e., xtabular* and mpxtabular* for use in two column mode where the table is meant to span both columns.

Table 1 was produced by code similar to the following:

\topcaption{The principal xtab package commands} \label{tab:xtab}
\tablefirsthead{\hline \multicolumn{1}{|c|}{\textbf{Command}} &
                       \multicolumn{1}{c|}{\textbf{Effect}} \\ \hline }
\tablehead{\multicolumn{2}{c}%
           {{\captionsize\bfseries \tablename\ \thetable{} --
             continued from previous page}} \\
  \hline   \multicolumn{1}{|c|}{\textbf{Command}} &
           \multicolumn{1}{c|}{\textbf{Effect}} \\ \hline }
\tablelasthead{\multicolumn{2}{c}%
           {{\captionsize\bfseries \tablename\ \thetable{} --
             concluded from previous page}} \\
  \hline   \multicolumn{1}{|c|}{\textbf{Command}} &
           \multicolumn{1}{c|}{\textbf{Effect}} \\ \hline }
\tabletail{\hline \multicolumn{2}{|r|}{{Continued on next page}} \\ \hline}
\tablelasttail{\hline \hline}
\begin{center}
\begin{xtabular}{|l|p{0.5\textwidth}|}
\verb|\begin{xtabular}{...}| & This is equivalent to the normal
                         \verb|\begin{tabular}{...}| environment.
                         You supply the specification of the columns
                         just as for the normal \Lenv{tabular} environment.
 \\
 &
                         All commands that can be used within a \Lenv{tabular}
                         environment can also be used within
                         the \Lenv{xtabular} environment.
 \\
 &
    Unlike the \Lenv{tabular} environment which prevents page breaking
within the tabular, the \Lenv{xtabular} allows page breaking, so that
tabulars can extend automatically across several pages.
... ... ...
\verb|\tablelasttail{...}| & The contents of this command are inserted before

                          the final (internal) \verb|\end{tabular}| of the table.
  \\
  &
     For example, you might want to note that  this is where
 the table ends.
  \\
 \end{xtabular}
 \end{center}

The table is only broken between rows — a row will not be split across pages. This can lead to some bad page breaks, especially if there are rows with a large vertical height (like some in Table 1). It is best to keep rows not too tall.

Unkike the table environment which floats, an xtabular environment is typeset at the point in the document where the environment is specified. It is best not to start an xtabular too close to the bottom of a page otherwise there might be an ugly page break.

\shrinkheight The command \shrinkheight{⟨length⟩} may be used after the first \\ in the table to modify the allowed height of the table on that page. A positive ⟨length⟩ decreases the allowed space on the page and a negative ⟨length⟩ increases the allowed space.

For example:

\shrinkheight{2\baselineskip} decreases the space per page by two lines. \shrinkheight{-\baselineskip} increases the space per page by one line. Note that I have never tried using this command so I cannot comment on its

efficacy. Instead, I use the \xentrystretch command when necessary.

\xentrystretch The command \xentrystretch{⟨decimal-fraction⟩} can be used before a table to modify the amount of vertical space apparently consumed by each entry in the subsequent table(s). The default is \xentrystretch{0.1} which specifies a 10% overestimate in the vertical space. Similarly, \xentrystretch{0.25} will overestimate the space by 25%. A different value may be used for each table in order to eliminate, or at least reduce, bad page breaks. Increasing the value causes fewer entries to be put on a page, thus reducing the chance of TEX putting in a page break before the xtab package is prepared for one.

You may specify the font used for the \tablehead and \tablelasthead yourself.

Note: Within ISO documents, captions shall be in bold font. The iso class also provides a command for setting the size of the font used in captions, namely \captionsize. The default value for this is set by the iso class. For the curious, the default definition is:

   \newcommand{\captionsize}{\normalsize}

2.1 Options

The xtab package has three options which control the amount of information that is written to the .log file. The options are:

1. The option errorshow (the default) does not write any extra information;

2. The option pageshow writes information about when and why xtab decides to produce a new page; 3. The option debugshow, which also includes pageshow, additionally writes

information about each line that is added to the table.

Under normal circumstances xtab is used without invoking any option. The pageshow option may be useful when attempting to cure a bad page break. The debugshow option, as its name implies, is principally of use to the xtab developer.

Independently of the options, the command \sstraceon may be used at any point in the document to turn on printing of debugshow data. This can be turned off later by the \sstraceoff command, which will stop all … show printing.

3 The implementation

The xtab package provides an extension to the supertabular package written by Johannes Braams and Theo Jurriens.^{3} The major portion of the following documentation is taken from supertabular.dtx. The package is designed to be used with the iso class in addition to the usual article, etc., classes.

The extension provided here enables the heading on the table on the last page to differ from those on earlier pages of the table. The implementation of the extension is based on ideas in David Carlisle’s longtable package. The downside of the extension is that L^{A}TEX has to be run twice if the document contains a supertabular. However, L^{A}TEX is usually run at least twice for any but the simplest document in order to get cross-references and Table of Contents, etc., resolved correctly.

The current version of the extension also either cures or reduces following weaknesses in the supertabular package.^{4}

1. Sometimes the top caption of a supertabular is printed on one page and the body is printed on the following page(s). That is, there is a lonely caption.

2. Sometimes the last page of a supertabular consists of an empty table. That is, just the head and foot of the table are printed.

3. If the number of lines in the first header for the table differs from the number of lines in subsequent headers, then the continuation pages of the table may be too short or, more troubling, too long.

The first version of xtab imported much of the code from the supertabular package (version 3.7) but I found that this did not work well because there were incompatible coded versions of supertabular available on CTAN. Further, I found that there were some problems with the original supertabular code in any case.^{5} I have to make the assumption that other users may have dissimilar or problematic versions, so include all the code here, and thus any errors can now be laid at my door.

The requirement for compatibilty with the iso class is achieved by modifications to the \ST@caption command only. Effectively this is orthogonal to the code required to implement the extension.

Now for the code itself. As syntactic sugar, all new macros for the extension have the prefix ‘PWST’ to distinguish them from the original macros. I have also denoted all extensions to the original supertabular by introducing them as Extension:.

3supertabular.sty, version 4.1c, 7 November 1997.

4Only the first two of these have been recognised by the authors of supertabular.

5I also found a bug in the 4.1b version which the authors kindly fixed in version 4.1c.

Announce the name and version of the package, which requires L^{A}TEX 2ε.

1 ⟨∗xtab⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{xtab}[2000/04/09 v2.3 Extended supertabular package]

4

\c@tracingst There are three options with the package which control the amount of information written to the log file:

1. errorshow (the default) no extra information

2. pageshow writes information about page breaking

3. debugshow adds information about each line that is added to the tabular

 5 \newcount\c@tracingst
 6 \DeclareOption{errorshow}{\c@tracingst\z@}

Extension: The next line in the original code did not do what the authors intended; the number should have been 3 rather than 2.

 7 %%%%\DeclareOption{pageshow}{\c@tracingst\tw@}
 8 \DeclareOption{pageshow}{\c@tracingst\thr@@}
 9 \DeclareOption{debugshow}{\c@tracingst5\relax}
10 \ProcessOptions

11

\if@topcaption \topcaption \bottomcaption The user-commands \topcaption and \bottomcaption set the flag @topcaption

to determine where to put the tablecaption. The default is to put the caption on the top of the table

12 \newif\if@topcaption \@topcaptiontrue
13 \def\topcaption{\@topcaptiontrue\tablecaption}
14 \def\bottomcaption{\@topcaptionfalse\tablecaption}

15

\PWST@thecaption \PWSTcapht Extension: \PWST@thecaption is used to store the text of the table’s caption. The vertical space required by a caption is stored in \PWSTcapht.

16 \gdef\PWST@thecaption{}
17 \newdimen\PWSTcapht

\tablecaption \@xtablecaption This command has to function exactly like \caption does except it has to store its argument (and the optional argument) for later processing within the supertabular environment.

18 \long\def\tablecaption{%
19 \refstepcounter{table}\@dblarg{\@xtablecaption}}
20 \long\def\@xtablecaption[#1]#2{%

Extension: I store the caption text for later measurement.

21 \long\gdef\PWST@thecaption{#2}%

Finish up with the original code.

22 \long\gdef\@process@tablecaption{\ST@caption{table}[#1]{#2}}}
23 \global\let\@process@tablecaption\relax

24

\ifST@star This switch is used in the internal macros to remember which kind of environment was started.

25 \newif\ifST@star

\ifST@mp This flag is used in the internal macros to remember if the tabular is to be put in a minipage.

26 \newif\ifST@mp

\ST@wd For the supertabular* environment it is necessary to store the intended width of the tabular.

27 \newdimen\ST@wd
\ST@rightskip \ST@leftskip \ST@parfillskip For the mpsupertabular environments we need special versions of \leftskip,
\rightskip and \parfillskip.
28 \newskip\ST@rightskip
29 \newskip\ST@leftskip
30 \newskip\ST@parfillskip

31

\@initisotab Required for ISO class, and check if class loaded.

32 \@ifundefined{@initisotab}{%
33 \newcommand{\@initisotab}{}
34 \newif\ifinfloat}{\typeout{xtab using iso captions}}

35

\ST@caption This is a redefinition of LaTeX’s \@caption, \@makecaption is called within a group so as not to return to \normalsize globally. In the original a fix was made for the ‘feature’ of the \@makecaption of article.sty and friends that a caption always gets a \vskip 10pt at the top and none at the bottom; if a user wants to precede his table with a caption this results in a collision. This fix is not implemented here as I think it should be done by the user modifying

\beforecaptionskip and \aftercaptionskip.

Extension: The ISO captioning is also initialised.

36 \long\def\ST@caption#1[#2]#3{\par%
37 \@initisotab
38 \addcontentsline{\csname ext@#1\endcsname}{#1}%
39  {\protect\numberline{%
40  \csname the#1\endcsname}{\ignorespaces #2}}
41 \begingroup
42  \@parboxrestore
43  \normalsize
44 %%  \if@topcaption \vskip -10\p@ \fi
45  \@makecaption{\csname fnum@#1\endcsname}{\ignorespaces #3}\par
46 %%  \if@topcaption \vskip 10\p@ \fi
47 \endgroup

Extension: The height of the caption is subtracted from the available space on the page.

48 \global\advance\ST@pageleft -\PWSTcapht
49 \ST@trace\tw@{Added caption. Space left for xtabular: \the\ST@pageleft}

50 }

51

\tablehead \tablefirsthead  \tablehead activates the new tabular \cr commands.
52 \newcommand\tablehead[1]{%
53 \gdef\@tablehead{%
54 \noalign{%
55  \global\let\@savcr=\\
56  \global\let\\=\org@tabularcr}%
57 #1%
58 \noalign{\global\let\\=\@savcr}}}
59 \tablehead{}
60 \newcommand\tablefirsthead[1]{\gdef\@table@first@head{#1}}

61

\c@PWSTtable \PWSTlastpage \PWSTcurpage \PWSTpenultimate \PWSTtempc \PWSTlines \PWSThead \PWSTlasthead Extension: These are counters for the supertabular extension. c@PWSTtable

counts the number of supertabulars in case one or more are not captioned. PW-STlastpage is a counter holding the number of pages that a supertabular uses and PWSTpenultimate is the penultimate page. PWSTcurpage counts the current number of supertabular pages processed. PWSTtempc is a scratch counter for page processing.

62 \newcounter{PWSTtable}
63 \newcount\PWSTlastpage
64 \newcount\PWSTpenultimate
65 \newcount\PWSTcurpage
66 \newcount\PWSTtempc

Extension: PWSTlines is used to count the number of supertabular entry lines on a page. Estimates of the number of lines in the normal table heading is held by PWSThead, and similarly PWSTlasthead is for the number of lines in the last heading.

67 \newcount\PWSTlines
68 %%% \newcount\PWSThead
69 %%% \newcount\PWSTlasthead

\iffirstcall Extension: This is used by the extension code to flag if the presumed last page overflows. If overflow occurs, then firstcall is set to false.

70 \newif\iffirstcall

\PWST@lastht \PWST@generalht \PWST@ht Extension: The estimated height of a table header and tail (i.e., the height of an empty table) for the last page of a supertabular is stored in \PWSTlastht. Similarly, the corresponding height of an empty table on a general page (neither the first nor the last) is stored in \PWSTgeneralht. \PWST@ht is a scratch variable.

71 \newdimen\PWST@lastht
72 \newdimen\PWST@generalht
73 \newdimen\PWST@ht

74

\tablelasthead \@table@last@head \notablelasthead Extension: \tablelasthead is the extension user command to specify the heading

for the last page of a supertabular. The command \notablelasthead switches off the last heading. This has to be used if a last headed table precedes one that does not have a special last head.

75 \newcommand{\tablelasthead}[1]{\gdef\@table@last@head{#1}}
76 \newcommand{\notablelasthead}{\let\@table@last@head\relax}

Now initialize these commands.

77 \tablelasthead{}
78 \notablelasthead

\tabletail \tablelasttail \tabletail is the user command to specify the appearance of the bottom of each tabular on a page. Special treatment is given to the end of the supertabular via

the \tablelasttail command.

If the user uses an extra amount of tabular-data (like \multicolumn) in \tabletail TEX starts looping because of the definition of \nextline. So make \\ act like just a \cr inside this tail to prevent the loop. Save and restore the value of \\

79 \newcommand\tabletail[1]{%
80 \gdef\@tabletail{%
81  \noalign{%
82  \global\let\@savcr=\\
83  \global\let\\=\org@tabularcr}%
84  #1%
85  \noalign{\global\let\\=\@savcr}}}
86 \tabletail{}
87 \newcommand\tablelasttail[1]{\gdef\@table@last@tail{#1}}
88 \tablelasttail{}

89

\sttraceon \sttraceoff The original supertabular included a tracing mechanism to follow the decisions supertabular made about page breaking. This is now also used as a debugging mechanism for the extension.

90 \newcommand\sttraceon{\c@tracingst5\relax}
91 \newcommand\sttraceoff{\c@tracingst\z@}

\ST@trace A macro that gets the trace message as its argument

92 \newcommand\ST@trace[2]{%
93 \ifnum\c@tracingst>#1\relax
94  \GenericWarning
95  {(xtab)\@spaces\@spaces}
96  {Package xtab: #2}%
97 \fi

98 }

99

\ST@pageleft This register holds the estimate of the amount of space left over on the current page. This is used in the decision when to start a new page.

100 \newdimen\ST@pageleft
\shrinkheight \setSTheight  \shrinkheight is a command to diminish the value of \ST@pageleft if necessary.

\setSTheight sets the value of \ST@pageleft if necessary.

101 \newcommand*\shrinkheight[1]{%
102 \noalign{\global\advance\ST@pageleft-#1\relax}}
103 \newcommand*\setSTheight[1]{%
104 \noalign{\global\ST@pageleft=#1\relax}}

\xentrystretch \PWST@xentrystretch Extension: Provide a user and internal command for fudging the estimated space taken by a table entry. Initialise to 10% increase.

105 \newcommand{\xentrystretch}[1]{\def\PWST@xentrystretch{#1}}
106 \xentrystretch{0.1}

107

\ST@headht \ST@tailht The register \ST@headht holds the height of the first head of a supertabular.

The register \ST@tailht holds the height of the tail.

108 \newdimen\ST@headht
109 \newdimen\ST@tailht

\ST@pagesofar Register \ST@pagesofar stores the estimate of the amount of the page already filled up.

110 \newdimen\ST@pagesofar

\ST@pboxht The measured (total) height of a parbox argument.

111 \newdimen\ST@pboxht

\ST@lineht \ST@stretchht \ST@prevht The estimated height of a normal line is stored in \ST@lineht. The register \ST@stretchht is used to store the difference between the normal line height and the line height when \arraystretch has a non-standard value. This is used in the case when p-box entries are added to the tabular. \ST@prevht stores the height of the previous line to use it as an estimate for the height of the next line. This is needed for a better estimate of when to break the tabular.

112 \newdimen\ST@lineht
113 \newdimen\ST@stretchht
114 \newdimen\ST@prevht

\ST@toadd When a tabular row is ended with \\[...] we need to temporarily store the optional argument in \ST@toadd.

115 \newdimen\ST@toadd

\ST@dimen A private scratch dimension register.

116 \newdimen\ST@dimen

\ST@pbox A box register to store the contents of a parbox.

117 \newbox\ST@pbox

118

\ST@tabularcr \ST@xtabularcr \ST@argtabularcr These are redefinitions of \@tabularcr and \@xtabularcr. This is needed to

include \ST@cr in the definition of \@xtabularcr.

All redefined macros have names that are similar to the original names, except with a leading ‘ST’.

119 \def\ST@tabularcr{%
120 {\ifnum0=`}\fi
121 \@ifstar{\ST@xtabularcr}{\ST@xtabularcr}}
122 \def\ST@xtabularcr{%
123 \@ifnextchar[%]
124  {\ST@argtabularcr}%
125  {\ifnum0=`{\fi}\cr\ST@cr}}
126 \def\ST@argtabularcr[#1]{%

127 \ifnum0=`{\fi}%
128 \ifdim #1>\z@
129  \unskip\ST@xargarraycr{#1}
130 \else
131  \ST@yargarraycr{#1}%
132 \fi}

\ST@xargarraycr \ST@yargarraycr In this case we need to copy the value of the optional argument of \\ in our private

 register \ST@toadd.
133 \def\ST@xargarraycr#1{%
134 \@tempdima #1\advance\@tempdima \dp \@arstrutbox
135 \vrule \@height\z@ \@depth\@tempdima \@width\z@ \cr
136 \noalign{\global\ST@toadd=#1}\ST@cr}

Here we need to insert \ST@cr

137 \def\ST@yargarraycr#1{%
138 \cr\noalign{\vskip #1\global\ST@toadd=#1}\ST@cr}

139

\ST@startpbox The macros that deal with parbox columns need to be redefined, because we need to know the size of the parbox.

140 \def\ST@startpbox#1{%

To achieve our goal we need to save the text in box.

141 \setbox\ST@pbox\vtop\bgroup\hsize#1\@arrayparboxrestore}
\ST@astartpbox supertabular version of \@astartpbox.
142 \def\ST@astartpbox#1{%
143 \bgroup\hsize#1%
144 \setbox\ST@pbox\vtop\bgroup\hsize#1\@arrayparboxrestore}
\ST@endpbox \ST@aendpbox supertabular versions of \@endpbox and \@aendpbox.
145 \def\ST@endpbox{%
146 \@finalstrut\@arstrutbox\par\egroup
147 \ST@dimen=\ht\ST@pbox
148 \advance\ST@dimen by \dp\ST@pbox
149 \ifnum\ST@pboxht<\ST@dimen
150  \global\ST@pboxht=\ST@dimen
151 \fi
152 \ST@dimen=\z@
153 \box\ST@pbox\hfil}
154 \def\ST@aendpbox{%
155 \@finalstrut\@arstrutbox\par\egroup
156 \ST@dimen=\ht\ST@pbox
157 \advance\ST@dimen by \dp\ST@pbox
158 \ifnum\ST@pboxht<\ST@dimen
159  \global\ST@pboxht=\ST@dimen
160 \fi
161 \ST@dimen=\z@
162 \unvbox\ST@pbox\egroup\hfil}

163

\estimate@lineht Estimates the height of normal line taking \arraystretch into account. Also computes the difference between a ‘normal’ line and a stretched one.

164 \def\estimate@lineht{%
165 \ST@lineht=\arraystretch \baselineskip
166 \global\advance\ST@lineht by 1\p@
167 \ST@stretchht\ST@lineht\advance\ST@stretchht-\baselineskip
168 \ifdim\ST@stretchht<\z@\ST@stretchht\z@\fi
169 \ST@trace\tw@{Basic line height: \the\ST@lineht\MessageBreak%
170  Arrayed line height: \the\ST@stretchht}%
171 \global\advance\ST@lineht \PWST@xentrystretch\ST@lineht
172 \ST@trace\tw@{Stretched line height: \the\ST@lineht}%

173 }

174

\@calfirstpageht Estimates the space left on the current page and decides whether the tabular can be started on this page or on a new page. Aspects of the original code are modified for the extension.

175 \def\@calfirstpageht{%
176 \ST@trace\tw@{Calculating height of xtabular on first page}%

The TEX register \pagetotal contains the height of the page sofar, the L^{A}TEX register \@colroom contains the height of the column.

177 \global\ST@pagesofar\pagetotal
178 \global\ST@pageleft\@colroom
179 \ST@trace\tw@{Height of previous text = \the\pagetotal; \MessageBreak
180  Height of column = \the\ST@pageleft}%

When we are in twocolumn mode TEX may still be collecting material for the first column although there seems to be no space left. In this case we have to check against two times \ST@pageleft.

181 \if@twocolumn
182  \ST@trace\tw@{two column mode}%
183  \if@firstcolumn
184  \ST@trace\tw@{First column}%
185  \ifnum\ST@pagesofar > \ST@pageleft
186  \global\ST@pageleft=2\ST@pageleft
187  \ifnum\ST@pagesofar > \ST@pageleft
188  \newpage\@calnextpageht
189  \ST@trace\tw@{starting new page}%
190  \else

In this case we’re in the second column, so we have to compensate for the material in the first column.

191  \ST@trace\tw@{Second column}%
192  \global\advance\ST@pageleft -\ST@pagesofar
193  \global\advance\ST@pageleft -\@colroom
194  \fi

When \ST@pagesofar is smaller than \ST@pageleft TEX is still collecting material for the first column, so we can start a new tabular environment like we do on a single column page.

195  \else
196  \global\advance\ST@pageleft by -\ST@pagesofar
197  \global\ST@pagesofar\z@
198  \fi
199  \else

When we end up here, TEX has already decided it had enough material for the first column and is building the second column.

200  \ST@trace\tw@{Second column}%
201  \ifnum\ST@pagesofar > \ST@pageleft
202  \ST@trace\tw@{starting new page}%
203  \newpage\@calnextpageht
204  \else
205  \global\advance\ST@pageleft by -\ST@pagesofar
206  \global\ST@pagesofar\z@
207  \fi
208  \fi
209 \else

In one column mode there is a simple decision.

210  \ST@trace\tw@{one column mode}%
211  \ifnum\ST@pagesofar > \ST@pageleft
212  \ST@trace\tw@{starting new page}%
213  \newpage\@calnextpageht

When we are not starting a new page subtract the size of the material already on it from the available space.

214  \else
215  \global\advance\ST@pageleft by -\ST@pagesofar
216  \global\ST@pagesofar\z@
217  \fi
218 \fi
219 \ST@trace\tw@{Available height: \the\ST@pageleft}%

Now we need to know the height of the head of the table. In order to measure this we typeset it in a normal tabular environment.

220 \ifx\@@tablehead\@empty
221  \ST@headht=\z@
222 \else
223  \setbox\@tempboxa=\vbox{\@arrayparboxrestore
224  \ST@restore
225  \expandafter\tabular\expandafter{\ST@tableformat}%
226  \@@tablehead\endtabular}%
227  \ST@headht=\ht\@tempboxa\advance\ST@headht\dp\@tempboxa
228 \fi
229 \ST@trace\tw@{Height of head: \the\ST@headht}%

To decide when to start a new page, we need to know the vertical size of the tail of the table.

230 \ifx\@tabletail\@empty
231  \ST@tailht=\z@
232 \else
233  \setbox\@tempboxa=\vbox{\@arrayparboxrestore
234  \ST@restore
235  \expandafter\tabular\expandafter{\ST@tableformat}
236  \@tabletail\endtabular}
237  \ST@tailht=\ht\@tempboxa\advance\ST@tailht\dp\@tempboxa
238 \fi

We add the average height of a line to this because when we decide to continue the tabular we need to have enough space left for one line and the tail.

239 \advance\ST@tailht by \ST@lineht
240 \ST@trace\tw@{Height of tail: \the\ST@tailht}%
241 \ST@trace\tw@{Maximum space for xtabular: \the\ST@pageleft}%

Now we decide whether we can continue on the current page or whether we need to start a new page. We assume that the minimum height of a tabular is the height of the head and tail and one line of data. If that doesn’t fit, start a new page.

242 \@tempdima\ST@headht
243 \advance\@tempdima\ST@lineht
244 \advance\@tempdima\ST@tailht

Extension: I also add the height of the caption to the required space. The amount to be added depends on whether it is a top or bottom caption. Allowance is also made for skips around the caption.

245 \if@topcaption
246  \setbox\@tempboxa=\vbox{\PWST@thecaption}
247  \PWSTcapht=\ht\@tempboxa\advance\PWSTcapht\dp\@tempboxa
248  \advance\PWSTcapht by 20\p@
249 \else
250  \PWSTcapht = 10\p@
251 \fi
252 \ST@trace\tw@{Caption height: \the\PWSTcapht}%
253 \advance\@tempdima\PWSTcapht

Continue with the original code.

254 \ST@trace\tw@{Minimum height of xtabular: \the\@tempdima}%
255 \ifnum\@tempdima>\ST@pageleft
256  \ST@trace\tw@{starting new page}%

Extension: The next line in the original code is \newpage\@calnextpageht. I need to start a new page, making allowance for the space required by the caption.

257  \newpage
258  \global\ST@pageleft\@colroom
259  \global\advance\ST@pageleft by -\PWSTcapht
260  \global\ST@pagesofar=\z@

Finish up with the original code.

261 \fi
262 }  % end \@calfirstpageheight

263

\@calnextpageht This calculates the maximum height of the tabular on all subsequent pages of the supertabular environment.

264 \def\@calnextpageht{%
265 \ST@trace\tw@{Calculating height of xtabular on next page}%
266 \global\ST@pageleft\@colroom
267 \global\ST@pagesofar=\z@
268 \ST@trace\tw@{Maximum space for xtabular: \the\ST@pageleft}%

269 }

270

\PWSTcalchtlines Extension: A macro to calculate the space required by an empty table and the number of lines in an empty table.

The appropriate heads and tails are typeset in a temporary box so we can measure them.

271 \newcommand{\PWSTcalchtlines}{%

Measure the lasttail.

272 \setbox\@tempboxa=\vbox{\@arrayparboxrestore
273  \ST@restore
274  \expandafter\tabular\expandafter{\ST@tableformat}%
275  \@table@last@tail\endtabular}%
276 \PWST@ht=\ht\@tempboxa\advance\PWST@ht\dp\@tempboxa
277 \global\PWST@lastht = \PWST@ht

And repeat for the lasthead.

278 \setbox\@tempboxa=\vbox{\@arrayparboxrestore
279  \ST@restore
280  \expandafter\tabular\expandafter{\ST@tableformat}%
281  \@table@last@head\endtabular}%
282 \PWST@ht = \ht\@tempboxa\advance\PWST@ht\dp\@tempboxa
283 \global\advance\PWST@lastht by \PWST@ht
284 \ST@trace\tw@{Height of empty xtabular on last page = \the\PWST@lastht}%

Now repeat pretty well all of the above for a general table (i.e., one that is not on the first page nor the designated last page).

First the tail.

285 \setbox\@tempboxa=\vbox{\@arrayparboxrestore
286  \ST@restore
287  \expandafter\tabular\expandafter{\ST@tableformat}%
288  \@tabletail\endtabular}%
289 \PWST@ht=\ht\@tempboxa\advance\PWST@ht\dp\@tempboxa
290 \global\PWST@generalht = \PWST@ht

And on to the general head.

291 \setbox\@tempboxa=\vbox{\@arrayparboxrestore
292  \ST@restore
293  \expandafter\tabular\expandafter{\ST@tableformat}%
294  \@tablehead\endtabular}%
295 \PWST@ht = \ht\@tempboxa\advance\PWST@ht\dp\@tempboxa
296 \global\advance\PWST@generalht by \PWST@ht
297 } % end \PWSTcalchtlines

298

\PWSTcalnextpageht Extension: From some experiments that I ran it appeared as though the supertabular package ignored the possibility that the space required for the table header and tail on pages after the first one might be different. If the subsequent head/tail combination were longer (i.e., took more vertical space) then the table could overflow the page. This is an attempt to fix this problem by calculating the actual minimum space required after the first page.

The calculations are similar to, but simpler, than those for \@calfirstpageht.

299 \newcommand{\PWSTcalnextpageht}{%
300 \ifnum\PWSTcurpage = \PWSTpenultimate
301  \ST@trace\tw@{Calculating height of xtabular on last page}%

We are on the penultimate page, so get the height of the last head/tail.

302  \PWST@ht=\PWST@lastht

Otherwise I need the general page data.

303 \else
304  \ST@trace\tw@{Calculating height of xtabular on next general page}%
305  \PWST@ht=\PWST@generalht
306 \fi

Having dealt with the two cases, I can now calculate the minimum space for a supertabular on the following page.

307 \global\ST@pageleft\@colroom
308 \global\advance\ST@pageleft -\PWST@ht
309 \global\ST@pagesofar=\z@
310 \ST@trace\tw@{Available space for xtabular: \the\ST@pageleft}%

311 }

312

\x@supertabular The various supertabular environments share a lot of code. Thus, to avoid needless

repetition, the shared code is defined in this macro.

This macro has been modified as part of the supertabular extension.

313 \def\x@supertabular{%

First save the original definition of \tabular and then make it point to \inner@tabular. This is done to enable supertabular cells to contain a tabular environment without getting unexpected results when the supertabular would be split across this inner tabular environment.

314 \let\org@tabular\tabular
315 \let\tabular\inner@tabular

The same has to be done for the tabular* environment. The coding is more verbose.

316 \expandafter\let
317  \csname org@tabular*\expandafter\endcsname
318  \csname tabular*\endcsname
319 \expandafter\let\csname tabular*\expandafter\endcsname
320  \csname inner@tabular*\endcsname

Extension: The original code printed out the top caption at this point. If there is too little space on the first page of the table, the tabular data is printed on the following page. If this is the case (and its not known yet whether it is), then the caption should also be printed on the following page.

321 %%%%  \if@topcaption \@process@tablecaption \fi

Back to the original code. Save the original definition of \\.

322 \global\let\@oldcr=\\

Save the current value of \baselineskip, as we need it in the calculation of the average height of a line.

323 \def\baslineskp{\baselineskip}%

We have to check whether array.sty was loaded, because some of the internal macros have different names.

324 \ifx\undefined\@classix

Save old \@tabularcr and insert the definition of \@stabularcr.

325  \let\org@tabularcr\@tabularcr
326  \let\@tabularcr\ST@tabularcr

Activate the new parbox algorithm.

327  \let\org@startpbox=\@startpbox
328  \let\org@endpbox=\@endpbox
329  \let\@@startpbox=\ST@startpbox
330  \let\@@endpbox=\ST@endpbox
331 \else

When array.sty was loaded things are a bit different.

332  \let\org@tabularcr\@arraycr
333  \let\@arraycr\ST@tabularcr
334  \let\org@startpbox=\@startpbox
335  \let\org@endpbox=\@endpbox
336  \let\@startpbox=\ST@astartpbox
337  \let\@endpbox=\ST@aendpbox
338 \fi

Check if the head of the table should be different for the first and subsequent pages.

339 \ifx\@table@first@head\undefined
340  \let\@@tablehead=\@tablehead
341 \else
342  \let\@@tablehead=\@table@first@head
343 \fi

The first part of a supertabular may be moved to the next page if it doesn’t fit on the current page. Subsequent parts can not be moved; therefore we will have to switch the definition of \ST@skippart around.

344 \let\ST@skippage\ST@skipfirstpart

Now we can estimate the average line height and the height of the first page of the supertabular.

345 \estimate@lineht
346 \@calfirstpageht

Extension: Call the macro to initialize the extension code for this table.

347 \PWSTinit

Extension: At this point I know, and have adjusted for, the page on which the first part of the table will be printed. It should now be safe to print the top caption, if any. Unfortunately, in spite of everthing, the TEX page breaking mechanism might still think that there is too little space left.

348 \if@topcaption \@process@tablecaption \fi
349 \noindent

Extension: Finally, subtract the space required by the header and the tail (as these don’t update the available space when output).

350 \global\advance\ST@pageleft -\ST@headht%
351 \ST@trace\tw@{Available space after accounting for header: \the\ST@pageleft}%
352 \global\advance\ST@pageleft -\ST@tailht%
353 \ST@trace\tw@{Available space after accounting for tail: \the\ST@pageleft}%

354 }

355

\PWSTinit Extension: This routine initialises the extension data.

356 \newcommand{\PWSTinit}{%

At the end of processing each supertabular (see later) the number of pages consumed by the supertabular is written to the .aux file. At the start of a supertabular, after incrementing the number of supertabulars processed, the prior number of pages are read from the file. These are stored in PWSTlastpage.

357 \global\advance\c@PWSTtable\@ne
358 \global\expandafter\let\expandafter\PWSTtempc
359  \csname PWST@\romannumeral\c@PWSTtable\endcsname

I have to take account of the fact that there might be no entry in the .aux file, and hence the lastpage number might not be set.

360 \ifx\PWSTtempc\relax
361  \ST@trace\tw@{Table \the\c@PWSTtable: Processing for the first time}%
362  \PWSTlastpage=\@m % = 1000
363 \else
364  \PWSTlastpage=\PWSTtempc
365 \fi
366 \ST@trace\tw@{Table \the\c@PWSTtable: last page set to \the\PWSTlastpage}%

Set the current page counter to unity.

367 \PWSTcurpage=\@ne

Perform the calculations for the empty table data.

368 \PWSTcalchtlines

Initialise the line counter and set firstcall to TRUE.

369 \global\PWSTlines=\z@
370 \global\firstcalltrue

If we have the iso class, then I have to flag that we are in a ‘float’.

371  \infloattrue

372 }

373

\xtabular We start by looking for an optional argument, which will be duly ignored as it

seems to make no sense to try to align a multipage table in the middle...

Extension: Use xtabular instead of supertabular, and similarly for the others, so this will not be mentioned explicitly again.

374 \def\xtabular{%
375 \@ifnextchar[{\@supertabular}%]
376  {\@supertabular[]}}

\@supertabular We can now save the preamble of the tabular in a macro.

377 \def\@supertabular[#1]#2{%
378 \def\ST@tableformat{#2}
379 \ST@trace\tw@{Starting a new xtabular}%

Then remember that this is not a supertabular* environment.

380 \global\ST@starfalse

Don’t use minipages.

381 \global\ST@mpfalse

Most of the following code is shared between the supertabbular and supertabular* environments. So to avoid duplication it is stored in a macro.

382 \x@supertabular

Finally start a normal tabular environment.

383 \expandafter\org@tabular\expandafter{\ST@tableformat}%
384 \@@tablehead}

385

\xtabular* We start by looking for the optional argument of the tabular environment.

386 \@namedef{xtabular*}#1{%
387 \@ifnextchar[{\@nameuse{@supertabular*}{#1}}%
388  {\@nameuse{@supertabular*}{#1}[]}%]

389 }

We start by saving the intended width and the preamble of the tabular*.

390 \@namedef{@supertabular*}#1[#2]#3{%
391 \ST@trace\tw@{Starting a new xtabular*}%
392 \def\ST@tableformat{#3}%
393 \ST@wd=#1\relax
394 \global\ST@startrue
395 \global\ST@mpfalse

Now we can call the common code for both environments.

396 \x@supertabular

And we can start a normal tabular* environment.

397 \expandafter\csname org@tabular*\expandafter\endcsname
398 \expandafter{\expandafter\ST@wd\expandafter}%
399 \expandafter{\ST@tableformat}%
400 \@@tablehead}

401

\mpxtabular This version of the supertabular environment puts each tabular into a minipage, thus making footnotes possible. We start by looking for an optional argument, which will be ignored as it makes no sense to try and align a multipage table in the middle…

402 \def\mpxtabular{%
403 \@ifnextchar[{\@mpsupertabular}%]
404  {\@mpsupertabular[]}}

We can now save the preamble in a macro.

405 \def\@mpsupertabular[#1]#2{%
406 \def\ST@tableformat{#2}%
407 \ST@trace\tw@{Starting a new mpxtabular}%

Remember that this is not a mpsupertabular* environment and also note we have to close the minipage later.

408 \global\ST@starfalse
409 \global\ST@mptrue

Since we are about to start a minipage of \columnwidth the horizontal alignment will not work. We have to remember the values and then restore them inside the minipage.

410 \ST@rightskip \rightskip
411 \ST@leftskip \leftskip
412 \ST@parfillskip \parfillskip

Call the code that is common to all the environments.

413 \x@supertabular

Finally, start a normal tabular

414 \minipage{\columnwidth}%
415 \parfillskip\ST@parfillskip
416 \rightskip \ST@rightskip
417 \leftskip \ST@leftskip
418 \noindent\expandafter\org@tabular\expandafter{\ST@tableformat}%
419 \@@tablehead}

420

\mpxtabular* We start by looking for the optional argument of the tabular environment.

421 \@namedef{mpxtabular*}#1{%
422 \@ifnextchar[{\@nameuse{@mpsupertabular*}{#1}}%
423  {\@nameuse{@mpsupertabular*}{#1}[]}%]

424 }

Now we can save the intended width and the preamble of the tabular*.

425 \@namedef{@mpsupertabular*}#1[#2]#3{%
426 \ST@trace\tw@{Starting a new mpxtabular*}%
427 \def\ST@tableformat{#3}%
428 \ST@wd=#1\relax
429 \global\ST@startrue
430 \global\ST@mptrue
431 \ST@rightskip \rightskip
432 \ST@leftskip \leftskip
433 \ST@parfillskip \parfillskip

Now is the time to call the common code for both environments.

434 \x@supertabular

And we can start a normal tabular* environment.

435 \minipage{\columnwidth}%
436 \parfillskip\ST@parfillskip
437 \rightskip \ST@rightskip
438 \leftskip \ST@leftskip
439 \noindent\expandafter\csname org@tabular*\expandafter\endcsname
440 \expandafter{\expandafter\ST@wd\expandafter}%
441 \expandafter{\ST@tableformat}%
442 \@@tablehead}%

443

\endxtabular \endxtabular* These close the xtabular and xtabular* environments.

For the extension, this macro has been modified to write out to the .aux file the number of pages used for the supertabular.

444 \def\endxtabular{%
445 \ifx\@table@last@tail\undefined
446  \@tabletail
447 \else
448  \@table@last@tail
449 \fi
450 \csname endtabular\ifST@star*\fi\endcsname

While studying the original code to determine where additions were needed for the extension, I realized that the last part of the \end... code was common to all the environments. I have broken it out into a seperate routine which also includes the modification needed for the extension.

451 \x@endsupertabular

And back to the original code.

452 \ST@trace\tw@{Ended a xtabular\ifST@star*\fi}%

453 }

The definition of the ending of the xtabular* environment is simple:

454 \expandafter\let\csname endxtabular*\endcsname\endxtabular

\x@endsupertabular This macro contains the code that is common to all the \end... commands. It includes the modification required for the extension.

455 \newcommand{\x@endsupertabular}{%

Restore the original definition of \@tabularcr

456 \ST@restore

Check if we have to insert a caption and restore to default behaviour of putting captions at the top.

457 \if@topcaption
458 \else
459  \@process@tablecaption
460  \global\@topcaptiontrue
461 \fi

Restore the meaning of \\ to the one it had before the start of this environment. Also re-initialize some control-sequences

462 \global\let\\=\@oldcr
463 \global\let\@table@first@head\undefined
464 %%%  \global\let\@table@last@tail\undefined
465 \global\let\@process@tablecaption\relax

Extension: For the extension, write the number of the last page to the .aux file. Also, if we are in the iso class, reset the ‘float’ flag.

466 \PWSToplastpagenum
467  \infloatfalse

468 }

469

\PWSToplastpagenum Extension: This routine is responsible for writing the number of the last page of

the supertabular to the .aux file.

What gets written is \PWST@vi{4}, assuming that the value of c@PWSTtable is 6 and the value of PWSTcurpage is 4.

470 \newcommand{\PWSToplastpagenum}{%

There are a number of cases to consider. The first decision is whether the current page is the previously calculated last page.

471 \ifnum\PWSTcurpage=\PWSTlastpage

The current table ends on the calculated last page. There are four cases to consider:

1. The table has not overflowed (firstcall is TRUE) and the table is not empty — this page is still the last page.

2. The table has not overflowed (firstcall is TRUE) and the table is empty — this page is after the actual last page, so decrease the page number.

3. The table has overflowed (firstcall is FALSE) and the overflow is large enough to generate a non-empty table on the next page — increment the page number.

4. The table has overflowed (firstcall is FALSE) and the overflow is small enough to generate an empty table on the next page — this page is still the last page.

472  \iffirstcall  % on last, no overflow
473 %%      \ifnum\PWSTlines < \PWSTlasthead % this table is empty
474  \ifnum\PWSTlines < \@ne            % this table is empty
475  \global\advance\PWSTcurpage \m@ne
476  \fi
477  \else % overflow
478 %%      \ifnum\PWSTlines > \tw@ % enough for another page
479  \ifnum\PWSTlines > \z@    % enough for another page
480  \global\advance\PWSTcurpage \@ne
481  \fi
482  \fi
483 \else

The table has ended on a page that is not the calculated last page. If the table is empty, then decrement the page number, else this is the last page.

484 %%    \ifnum\PWSTlines < \PWSThead % empty table
485  \ifnum\PWSTlines < \@ne        % empty table
486  \global\advance\PWSTcurpage \m@ne
487  \fi
488 \fi

Finally, write out the ‘new’ last page number.

489 \if@filesw\immediate\write\@auxout%
490  {\gdef\string\PWST@\romannumeral\c@PWSTtable{\the\PWSTcurpage}}
491  \ST@trace\tw@{Table \the\c@PWSTtable:\MessageBreak
492  wrote \the\PWSTcurpage\space as the last page}%
493 \fi

494 }

495

\endmpxtabular \endmpxtabular* These close the mpxtabular and mpxtabular* environments.
496 \def\endmpxtabular{%
497 \ifx\@table@last@tail\undefined
498  \@tabletail
499 \else
500  \@table@last@tail
501 \fi
502 \csname endtabular\ifST@star*\fi\endcsname
503 \endminipage

Now call the common code for all \end....

504 \x@endsupertabular

Finish per the original code.

505 \ST@trace\tw@{Ended an mpxtabular\ifST@star*\fi}%

506 }

The definition of the ending of the mpxtabular* environment is simple:

507 \expandafter\let\csname endmpxtabular*\endcsname\endmpxtabular

508

\ST@restore This macro restores the original definitions of the macros that handle parbox entries and the ‘end of row’ macros.

509 \def\ST@restore{%
510 \ifx\undefined\@classix
511  \let\@tabularcr\org@tabularcr
512 \else
513  \let\@arraycr\org@tabularcr
514 \fi
515 \let\@startpbox\org@startpbox
516 \let\@endpbox\org@endpbox

517 }

518

\inner@tabular \inner@tabular* In order to facilitate complete tabular environments to be in a cell of a supertabular we need to adapt the definition of the original environments. For the inner tabular a number of definitions have to be restored.

519 \def\inner@tabular{%
520 \ST@restore
521 \let\\=\@oldcr
522 \noindent
523 \org@tabular}
524 \@namedef{inner@tabular*}{%
525 \ST@restore
526 \let\\=\@oldcr
527 \noindent
528 \csname org@tabular*\endcsname}

529

\ST@cr This macro is called by each \\ inside the tabular environment. It updates the estimate of the amount of space left on the current page and starts a new page if necessary.

530 \def\ST@cr{%
531 \noalign{%
532  \ST@trace\thr@@{Parbox height: \the\ST@pboxht\MessageBreak
533  Line height: \the\ST@lineht}%
534  \ifnum\ST@pboxht<\ST@lineht

If there is a non-empty line, but an empty parbox, then \ST@pboxht might be non-zero, but too small thereby breaking the algorithm. Therefore we estimate the height of the line to be \ST@lineht in this case, and store it in \ST@prevht.

535  \global\advance\ST@pageleft -\ST@lineht
536  \global\ST@prevht\ST@lineht
537  \else

When the parbox is not empty we take its height into account plus a little extra.

538  \global\advance\ST@pboxht \PWST@xentrystretch\ST@pboxht
539  \global\advance\ST@pboxht \ST@stretchht
540  \ST@trace\thr@@{Added par box with height \the\ST@pboxht}%
541  \global\advance\ST@pageleft -\ST@pboxht
542  \global\ST@prevht\ST@pboxht

543  \global\ST@pboxht\z@
544  \fi

\ST@toadd is the value of the optional argument of \\.

545  \global\advance\ST@pageleft -\ST@toadd
546  \global\ST@toadd=\z@
547  \ST@trace\thr@@{Space left for xtabular: \the\ST@pageleft}%

Extension: Increment the line number at this point.

548  \global\advance\PWSTlines \@ne
549  \ST@trace\thr@@{Line counter incremented by one to: \the\PWSTlines}%
550 } % end of noalign

In general, when the \ST@pageleft has become negative, the last row was so high that the supertabular doesn’t fit on the current page. In this case we skip the current page and start at the top of the next one; otherwise TEX will move this part of the table to a new page anyway, probably with a message about an overfull

\vbox.

Extension: For the extension I do some special handling if we are on the last page. Essentially the idea is not to start a new page, but to continue on the current page, noting any overflow.

551 \ifnum\PWSTcurpage=\PWSTlastpage
552  \PWST@lastpagecr
553 \else

Execute the original code.

554  \ifnum\ST@pageleft<\z@
555  \ST@skippage
556  \else

When there is not enough space left on the current page, we start a new page. To compute the amount of space needed we use the height of the previous line (\ST@prevht) as an estimate of the height of the next line. If we are processing an mpsupertabular we also need to take the height of the footnotes into account.

557  \noalign{\global\@tempdima\ST@tailht
558  \global\advance\@tempdima\ST@prevht
559  \ifST@mp
560  \ifvoid\@mpfootins\else
561  \global\advance\@tempdima\ht\@mpfootins
562  \global\advance\@tempdima 3pt
563  \fi
564  \fi} % end noalign
565  \ifnum\ST@pageleft<\@tempdima
566  \ST@newpage
567  \else

This line is necessary because the tablehead has to be inserted after the \if\else\fi-clause. For this purpose \ST@next is used. In the middle of tableprocessing it should be an empty macro (not \relax).

568  \noalign{\global\let\ST@next\@empty}%
569  \fi
570  \fi

Extension: Close off the \iflastpage;

571 \fi

and finish per the original code.

572 \ST@next}

573

\PWST@lastpagecr Extension: This routine handles newlines on the last page of a supertabular. The idea is that when we are on the last page the table continues to be processed until the end without calling for a newpage even if the table will be too long. I do need to record whether or not the table has ‘overflowed’ the allowable space on the page. The code is very similar to the last part of the code for \ST@cr.

574 \newcommand{\PWST@lastpagecr}{%
575 \noalign{%
576  \ifnum\ST@pageleft<\z@

The table has overflowed, so record the fact.

577  \PWST@setfirstcall
578  \fi

Now continue along the lines of \ST@cr.

579  \global\@tempdima\ST@tailht
580  \global\advance\@tempdima\ST@prevht
581  \ifST@mp
582  \ifvoid\@mpfootins\else
583  \global\advance\@tempdima\ht\@mpfootins
584  \global\advance\@tempdima 3pt
585  \fi
586  \fi
587  \ifnum\ST@pageleft<\@tempdima

Again, the table has overflowed.

588  \PWST@setfirstcall
589  \fi

Finish like \ST@cr.

590  \global\let\ST@next\@empty
591  } % end noalign

592 }

593

\PWST@setfirstcall Extension: This routine records that a table on the last page has overflowed by setting firstcall to FALSE. If it is the first such overflow it also zeroes the line counter.

594 \newcommand{\PWST@setfirstcall}{%
595 \iffirstcall
596  \global\firstcallfalse
597  \global\PWSTlines=\z@
598  \ST@trace\thr@@{Overflow on last page. Line counter set to \the\PWSTlines}%
599 \fi

600 }

601

\ST@skipfirstpart This macro skips the current page and moves the entire supertabular that has been built so far to the next page.

602 \def\ST@skipfirstpart{%
603 \noalign{%
604  \ST@trace\tw@{Tabular too high, moving to next page}%

In order for this to work properly we need to adapt the value of \ST@pageleft. When this macro is called it has a negative value. We should add the height of the next page to that (\@colroom). From the result the ‘normal’ height of the supertabular should be subtracted (\@colroom - \pagetotal). This could be coded as follows:

      \ST@dimen\@colroom
      \advance\ST@dimen-\pagetotal
      \global\advance\ST@pageleft\@colroom
      \global\advance\ST@pageleft-\ST@dimen

However, note that \@colroom is added and subtracted. Thus the code can be simplified to:

605  \global\advance\ST@pageleft\pagetotal

Then we can set \ST@pagesofar to zero and start the new page.

606  \global\ST@pagesofar\z@
607  \newpage

Finally we make sure that this macro can only be executed once for each supertabular by changing the definition of \ST@skippage.

608  \global\let\ST@skippage\ST@newpage

609 } 610 }

611

\ST@newpage This macro performs the actions necessary to start a new page. This macro is also modified for the extension to supertabluar.

612 \def\ST@newpage{%
613 \noalign{\ST@trace\tw@{Starting new page, writing tail}}%

Output \tabletail, close the tabular environment, close a minipage if necessary, output all material and start a fresh new page.

614 \@tabletail
615 \ifST@star
616  \csname endtabular*\endcsname
617 \else
618  \endtabular
619 \fi
620 \ifST@mp
621  \endminipage
622 \fi

Then we make sure that \ST@skippage can no longer be executed for this supertabular by changing its definition.

623 \global\let\ST@skippage\ST@newpage

On with the output.

Extension: The original code had the next line as \newpage\@calnextpageht. However, if the general header has a vertical height that differs from the first header, then the table on the continuation pages may run short or, more disconcerting, long. The extension, I think, cures that by using a different algorithm to calculate the height on the next page.

624 \newpage\PWSTcalnextpageht
625 \ST@trace\tw@{writing head}%

Extension: The original code just let \ST@next to \@tablehead. The extension has to handle the special case of of the heading on the last page.

626 \PWSTsethead

Now we are back to the original supertabular code.

627 \ifST@mp
628  \noindent\minipage{\columnwidth}%
629  \parfillskip\ST@parfillskip
630  \rightskip \ST@rightskip
631  \leftskip \ST@leftskip
632 \fi
633 \noindent
634 \ifST@star
635  \expandafter\csname org@tabular*\expandafter\endcsname
636  \expandafter{\expandafter\ST@wd\expandafter}%
637  \expandafter{\ST@tableformat}%
638 \else
639  \expandafter\org@tabular\expandafter{\ST@tableformat}%
640 \fi}

641

\PWSTsethead Extension: This is more extension code for use within \ST@newpage. It provides the proper table head for the page about to be processed.

642 \newcommand{\PWSTsethead}{%

First the line counter is zeroed.

643 \global\PWSTlines=\z@
644 \ST@trace\thr@@{Newpage, line counter set to: \the\PWSTlines}%

The current page counter is incremented and it is checked against the old page counter to see if this is the last page of this supertabular.

645 \global\advance\PWSTcurpage\@ne
646 \ST@trace\tw@{Table \the\c@PWSTtable:\MessageBreak
647  current page = \the\PWSTcurpage,\MessageBreak
648  last page = \the\PWSTlastpage}%
649 \ifnum\PWSTcurpage=\PWSTlastpage
650  \ST@trace\tw@{Newpage is the last page}%

We are on the last page. If there are more than one pages and the last table heading has been specified, then the heading is set to \@table@last@head, otherwise it is

 set to \@tablehead.
651  \ifnum\PWSTcurpage>\@ne
652  \ifx\@table@last@head\relax
653  \let\ST@next\@tablehead
654  \ST@trace\tw@{Set heading to tablehead}%
655  \else
656  \let\ST@next\@table@last@head
657  \ST@trace\tw@{Set heading to tablelasthead}%
658  \fi
659  \fi
660 \else

We are not on the last page, so just set the heading to \@tablehead.

661  \let\ST@next\@tablehead
662  \ST@trace\tw@{Set heading to tablehead}%

663 \fi

664 }

665

The end of this package

666 ⟨/xtab⟩

References

[GMS94] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The LaTeX Companion. Addison-Wesley Publishing Company, 1994.

[Wil96] Peter R. Wilson. LaTeX for standards: The LaTeX package files user manual. NIST Report NISTIR, June 1996.

Index

Numbers written in italic refer to the page where the corresponding entry is described; numbers underlined refer to the code line of the definition; numbers in roman refer to the code lines where the entry is used.

Symbols

\@@endpbox . . . . . . . 330
\@@startpbox . . . . . . 329
\@@tablehead . . 220,

226, 340, 342, 384, 400, 419, 442

\@arraycr . 332, 333, 513

\@arrayparboxrestore

. . . . . . . . 141, 144, 223, 233, 272, 278, 285, 291

\@arstrutbox . . . . . .

…. 134, 146, 155 \@auxout . . . . . . . . . 489

\@calfirstpageheight

. . . . . . . . . . . 262

\@calfirstpageht . .

. . . . . . . . \underline{175}, 346

\@calnextpageht . . .

. 188, 203, 213, \underline{264} \@classix …. 324, 510

\@colroom . . . . 178,

193, 258, 266, 307 \@dblarg . . . . . . . . . 19 \@depth . . . . . . . . . . 135

\@endpbox . . . . . . . .

. 328, 335, 337, 516

\@finalstrut . . 146, 155

\@height . . . . . . . . . 135

\@ifnextchar . . 123,

375, 387, 403, 422 \@ifstar . . . . . . . . . 121

\@ifundefined . . . . . 32
\@initisotab . . . . \underline{32}, 37
\@makecaption . . . . . 45
\@mpfootins . . . . . . .

. 560, 561, 582, 583

\@mpsupertabular . .

. . . . . . . . 403–405

\@namedef . . . . 386,

390, 421, 425, 524

\@nameuse . . . . . . . .

. 387, 388, 422, 423 \@oldcr 322, 462, 521, 526

\@parboxrestore . . . 42
\@process@tablecaption

. . . . . . 22, 23, 321, 348, 459, 465

\@savcr . . 55, 58, 82, 85 \@spaces . . . . . . . . . 95

\@startpbox . . . . . . .

. 327, 334, 336, 515

\@supertabular . . . .

…. 375, 376, \underline{377}

\@table@first@head .

. 60, 339, 342, 463

\@table@last@head .

. \underline{75}, 281, 652, 656

\@table@last@tail .

. . 87, 275, 445, 448, 464, 497, 500

                          \@tablehead . . . . 53,

294, 340, 653, 661

                          \@tabletail . . . . . . .

. . 80, 230, 236, 288, 446, 498, 614

                          \@tabularcr 325, 326, 511
                          \@topcaptionfalse . 14
                          \@topcaptiontrue . .

. . . . . . 12, 13, 460 \@width . . . . . . . . . . 135

                          \@xtablecaption . . . \underline{18}

A

                          \addcontentsline . . 38
                          \arraystretch . . . . . 165

B

                          \baselineskip . . . . .

…. 165, 167, 323

                          \baslineskp . . . . . . . 323
                          \bottomcaption . . . . \underline{12}

C

                          \c@PWSTtable . . . \underline{62},

357, 359, 361, 366, 490, 491, 646

                          \c@tracingst \underline{5}, 90, 91, 93
                          \columnwidth . . . . . .

…. 414, 435, 628 \csname . . . . . . . 38,

40, 45, 317–320, 359, 397, 439, 450, 454, 502, 507, 528, 616, 635

D

\dp . . 134, 148, 157, 227, 237, 247, 276, 282, 289, 295

E

\endcsname . . . . 38,

40, 45, 317–320, 359, 397, 439, 450, 454, 502, 507, 528, 616, 635

\endminipage . . 503, 621
\endmpxtabular . . . . \underline{496}
\endmpxtabular* . . . \underline{496}
\endtabular . . . . . . .

. 226, 236, 275, 281, 288, 294, 618

\endxtabular . . . . . . \underline{444}
\endxtabular* . . . . . \underline{444}
\estimate@lineht . .

. . . . . . . . \underline{164}, 345

F

\firstcallfalse . . . 596
\firstcalltrue . . . . 370

G

\GenericWarning . . . 94

H

\ht 147, 156, 227, 237, 247, 276, 282, 289, 295, 561, 583

I

\if@filesw . . . . . . . 489
\if@firstcolumn . . . 183
\if@topcaption . . . .

… \underline{12}, 44, 46, 245, 321, 348, 457

\if@twocolumn . . . . . 181
\iffirstcall \underline{70}, 472, 595
\ifinfloat . . . . . . . 34

\ifST@mp . . . . . . \underline{26},

559, 581, 620, 627

\ifST@star . . . . . . .

. . \underline{25}, 450, 452, 502, 505, 615, 634

\ifvoid . . . . . . 560, 582

\ignorespaces . . . 40, 45
\immediate . . . . . . . 489
\infloatfalse . . . . . 467
\infloattrue . . . . . . 371
\inner@tabular 315, \underline{519}
\inner@tabular* . . . \underline{519}

L

\leftskip . . . . 411,

417, 432, 438, 631

M

\MessageBreak . . . . .

…. 169, 179, 491, 532, 646, 647

\minipage . 414, 435, 628

\mpxtabular . . . . . . . \underline{402}
\mpxtabular* . . . . . . \underline{421}

N

\newbox . . . . . . . . . . 117 \newcount …. 5, 63–69 \newif . 12, 25, 26, 34, 70 \newskip . . . . . . . 28–30 \noalign 54, 58, 81, 85,

102, 104, 136, 138, 531, 557, 568, 575, 603, 613

\normalsize . . . . . . . 43
\notablelasthead . . \underline{75}

O

\org@endpbox . . . . . .

…. 328, 335, 516

\org@startpbox . . . .

…. 327, 334, 515

\org@tabular . . 314,

383, 418, 523, 639

\org@tabularcr 56, 83,

325, 332, 511, 513

P

\pagetotal 177, 179, 605

\ProvidesPackage . . . 3

\PWST@ . . . . . . . . . . . 490

\PWST@generalht . . .

. \underline{71}, 290, 296, 305 \PWST@ht … \underline{71}, 276,

277, 282, 283, 289, 290, 295, 296, 302, 305, 308

\PWST@lastht . . . \underline{71},

277, 283, 284, 302

\PWST@lastpagecr . .

. . . . . . . . 552, \underline{574}

                          \PWST@setfirstcall .

…. 577, 588, \underline{594}

                          \PWST@thecaption . .

. . . . . . \underline{16}, 21, 246

                          \PWST@xentrystretch

…. \underline{105}, 171, 538

                          \PWSTcalchtlines . .

. . . . . . . . \underline{271}, 368

                          \PWSTcalnextpageht .

. . . . . . . . \underline{299}, 624

                          \PWSTcapht . . . . . . .

\underline{16}, 48, 247, 248, 250, 252, 253, 259

                          \PWSTcurpage . . . \underline{62},

300, 367, 471, 475, 480, 486, 490, 492, 551, 645, 647, 649, 651

\PWSThead . . . . . \underline{62}, 484 \PWSTinit …. 347, \underline{356}

                          \PWSTlasthead . . \underline{62}, 473

                          \PWSTlastpage . . \underline{62},

362, 364, 366, 471, 551, 648, 649

                          \PWSTlines . . . . \underline{62},

369, 473, 474, 478, 479, 484, 485, 548, 549, 597, 598, 643, 644

                          \PWSToplastpagenum .

. . . . . . . . 466, \underline{470}

                          \PWSTpenultimate  \underline{62}, 300
                          \PWSTsethead . . 626, \underline{642}
                          \PWSTtempc . . . . . . .

. \underline{62}, 358, 360, 364

R

                          \romannumeral . 359, 490

S

                          \setSTheight . . . . . . \underline{101}
                          \shrinkheight . . . 6, \underline{101}
                          \ST@aendpbox . . \underline{145}, 337
                          \ST@argtabularcr . . \underline{119}
                          \ST@astartpbox   \underline{142}, 336
                          \ST@caption . . . . . 22, \underline{36}

\ST@cr 125, 136, 138, \underline{530}

                          \ST@dimen . . . . . . . .

. \underline{116}, 147–150, 152, 156–159, 161

                          \ST@endpbox . . . \underline{145}, 330

                          \ST@headht   \underline{108}, 221,

227, 229, 242, 350

                          \ST@leftskip  \underline{28}, 411,

417, 432, 438, 631

\ST@lineht . . . . . . .

. \underline{112}, 165–167, 169, 171, 172, 239, 243, 533–536

\ST@mpfalse . . . 381, 395
\ST@mptrue . . . 409, 430
\ST@newpage 566, 608, \underline{612}

\ST@next . . 568, 572,

590, 653, 656, 661

\ST@pageleft . . . . . .

… 48, 49, \underline{100}, 102, 104, 178, 180, 185–187, 192, 193, 196, 201, 205, 211, 215, 219, 241, 255, 258, 259, 266, 268, 307, 308, 310, 350– 353, 535, 541, 545, 547, 554, 565, 576, 587, 605

\ST@pagesofar . \underline{110},

177, 185, 187, 192, 196, 197, 201, 205, 206, 211, 215, 216, 260, 267, 309, 606

\ST@parfillskip . . .

. . . . . \underline{28}, 412, 415, 433, 436, 629

\ST@pbox . . \underline{117}, 141, 144, 147, 148, 153, 156, 157, 162

\ST@pboxht   \underline{111}, 149,

150, 158, 159, 532, 534, 538–543

\ST@prevht . . . \underline{112},

536, 542, 558, 580

\ST@restore . . . . . . .

. 224, 234, 273, 279, 286, 292, 456, \underline{509}, 520, 525

                   \ST@rightskip \underline{28}, 410,

416, 431, 437, 630

                   \ST@skipfirstpart .

. . . . . . . . 344, \underline{602}

                   \ST@skippage . . . . . .

. 344, 555, 608, 623

                   \ST@starfalse . 380, 408
                   \ST@startpbox . \underline{140}, 329
                   \ST@startrue . . 394, 429
                   \ST@stretchht . \underline{112},

167, 168, 170, 539

                   \ST@tableformat . . .

. 225, 235, 274, 280, 287, 293, 378, 383, 392, 399, 406, 418, 427, 441, 637, 639

                   \ST@tabularcr . . . . .

…. \underline{119}, 326, 333

                   \ST@tailht   \underline{108}, 231,

237, 239, 240, 244, 352, 557, 579

                   \ST@toadd . . . . \underline{115},

136, 138, 545, 546 \ST@trace 49, \underline{92}, 169,

172, 176, 179, 182, 184, 189, 191, 200, 202, 210, 212, 219, 229, 240, 241, 252, 254, 256, 265, 268, 284, 301, 304, 310, 351, 353, 361, 366, 379, 391, 407, 426, 452, 491, 505, 532, 540, 547, 549, 598, 604, 613, 625, 644, 646, 650, 654, 657, 662

\ST@wd . . . . . \underline{27}, 393, 398, 428, 440, 636

                                             \ST@xargarraycr 129, \underline{133}
                                             \ST@xtabularcr . . . . \underline{119}
                                             \ST@yargarraycr 131, \underline{133}

\string . . . . . . . . . . 490

                                             \sttraceoff . . . . . . . \underline{90}
                                             \sttraceon . . . . . . . \underline{90}

T

                                             \tablecaption 13, 14, \underline{18}
                                             \tablefirsthead . . . \underline{52}
                                             \tablehead . . . . . . . \underline{52}
                                             \tablelasthead . . . . \underline{75}
                                             \tablelasttail . . . . \underline{79}
                                             \tabletail . . . . . . . \underline{79}

\tabular . . . . . 225, 235, 274, 280, 287, 293, 314, 315

\thr@@ . . 8, 532, 540,

547, 549, 598, 644

                                             \topcaption . . . . . . . \underline{12}

U \unskip . . . . . . . . . . 129 \unvbox . . . . . . . . . . 162

V \vbox . 223, 233, 246,

272, 278, 285, 291 \vrule . . . . . . . . . . . 135 \vtop . . . . . . . . 141, 144

W \write . . . . . . . . . . . 489

X

                                             \x@endsupertabular .

…. 451, \underline{455}, 504

                                             \x@supertabular  \underline{313},

382, 396, 413, 434

                                             \xentrystretch . . 6, \underline{105}

\xtabular . . . . . . . . \underline{374}

                                             \xtabular* . . . . . . . \underline{386}